Howard Industries Inc., operating at full capacity, sold 64,000 units at a price of $45 per unit during the current year. Its income statement is as follows $2,880,000 1,400,000 $1,480,000 Expenses 400,000 387,500 787,500 692,500 Income from operations The division of costs between variable and fixed is as follows Cost of goods sold Selling expenses Administrative expenses Variable 75% 60% 80% Fixed 25% 40% 20% Management is considering a plant expansion program for the following year that will permit an increase of $900,000 in yearly sales. The expansion will increase fixed costs by $212,500 but wil not affect the relationship between sales and variable costsInstructions 1. Determine the total fixed costs and the total variable costs for the current year. 2. Determine (a) the unit variable cost and (b) the unit contribution margin for the current year. 3. Compute the break-even sales (units) for the current year. 4 Compute the break-even sales (units) under the proposed program for the 5. Determine the amount of sales (units) that would be necessary under the following year proposed program to realize the $692,500 of income from operations that was 6. Determine the maximum income from operations possible with the expanded 7. If the proposal is accepted and sales remain at the current level, what will the 8. Based on the data given, would you recommend accepting the proposal? Explain earned in the current year plant. income or loss from operations be for the following year?
Answer #1

Solution:
1)

Total variable costs   1,600,000     
Total fixed costs   587,500     
Working:

   Variable   Fixed     
COGS   1,050,000   350,000     
Selling exp   240,000   160,000     
Administrative Expense   310,000   77,500     
   1,600,000   587,500     

2)

Unit variable cost   $25     
Unit contribution margin   $20     

Working:
Variable cost per unit : 1,600,000 / 64,000 units = $25
CM = Selling price - VC per unit = 45-25 = $20
3) Break-even sales (units): 29,375 units
Working:
Break-even sales (units): Fixed cost / CM = 587,500 / 20 = 29,375
4) Break-even sales (units) under the proposed program: 40,000 units
Working:
Break-even sales (units): 587,500 + 212500 / 20 = 40,000 units